OpenStudy (anonymous):

Solve the equation -3 cos t= 1 in the interval from 0 < theta < 2pi. Round to the nearest hundredth.

zepdrix (zepdrix):

Hey :) \(\large\rm -3\cos t=1\) Dividing by -3 gives us: \(\large\rm \cos t=-\dfrac{1}{3}\)

zepdrix (zepdrix):

To solve for t, I suppose we need to make use of our inverse function, \(\large\rm \arccos\left(-\dfrac{1}{3}\right)=t\)